OpenStudy (anonymous):

Could someone help me solve this substitution problem? Would really appreciate it! The answer should be a coordinate. 4x = 3y - 2 18 = 3x + y Also, another one: 3x + 2y = 23 1/2x - 4 = y

jigglypuff314 (jigglypuff314):

\[\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange} {\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange} {\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\ \Huge\sf\color{blue}{~~~~Welcome ~To~ OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color {blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue} {\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue} {\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\] to solve by substitution for 4x = 3y - 2 18 = 3x + y --> y = -3x + 18 then you can plug that into the other equation 4x = 3y - 2 --> 4x = 3(-3x + 18) - 2 then solve for x (use distributive property etc...) then once you get x, plug it back into one of the original equations then solve for y.

OpenStudy (anonymous):

Thank you so much! I was mainly more confused on the second one, but this definitely helped a lot! I think it was the fraction in the second that that threw me off slightly. Would you be able to help me through the steps of that one as well? Would really appreciate it!

OpenStudy (anonymous):

Also, thank you for the welcome! I heard about this website just recently and thought I'd give it a try.

jigglypuff314 (jigglypuff314):

I got answers, so it's not impossible... but they're weird answers... 3x + 2y = 23 1/2x - 4 = y <--- substitute this one into the above equation so 3x + 2y = 23 --> 3x + 2(1/2 x - 4) = 23 distribute so 3x + 2(1/2)x - 2(4) = 23 then solve for x 3x + x - 8 = 23 add like terms and add 8 to both sides 4x = 31 then divide both sides by 4 x = 31/4 then once you get x, plug it back into one of the original equations 1/2x - 4 = y --> 1/2 (31/4) - 4 = y turned 4 into fraction to subtract easier 31/8 - 32/8 = y if it look better this way \[\frac{ 31 }{ 8 } - \frac{ 32 }{ 8 } = y\] so y = - 1/8 therefore the (x, y) coordinate would be (31/4, -1/8)